OpenStudy (anonymous):

Write the conic section in standard form, identify it, and graph and label the appropriate parts.

OpenStudy (anonymous):

\[3x ^{2}-5x + 3y ^{2}-8=0\]

OpenStudy (anonymous):

this is an ellipse

OpenStudy (anonymous):

how do you put it in standard form?

OpenStudy (anonymous):

first write the equation as x^2 -5x/3+y^2=8/3 now for factorising the x term , add and subtract 25/36....u'll end up gettin, (x-5/6)^2+y^2=121/36 its a circle(oops!) and centred about (5/6,0) with radius 11/6. ps:a circle is an ellipse too , when a=b.